A water treatment plant (10 MGD) uses 50 kg/d of chlorine. The residual chlorine after 30 minutes contact time is 0.55 mg/L. Find the chlorine dosage and chlorine demand of the water.
Answer #1

Amount of chlorine added (kg)/d = Q(MLD) * (mgly sokold = 37.854 * C. Ca 1.32 mg/h chlorine dosage (c): 1-32 mg/L. chlorine d